HUBBLE SPACE TELESCOPE DAILY REPORT #1353 PERIOD COVERED: 0000Z (UTC) 03/22/95 - 0000Z (UTC) 03/23/95 Daily Status Report as of 082/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 5433 (Testing the AGN Paradigm in Nearby Seyfert Galaxies: Cycle 4 High) The Faint Object Spectrograph (Red Side) was used to make spectral observations of galaxy NGC-3031-NU. The spectra was obtained at several positions both on and slightly off the nucleus of the galaxy. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.2 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.3 Completed WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-5457 to be used in measuring the periods of Cepheid variable stars within the galaxy. The observations were completed as planned, and no anomalies were noted. 1.4 Completed FOS/RD 5532 (FOS Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to monitor the instrument's background (dark) images. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5766 (Determination of the Extragalactic distance Scale - M-101 Carryover) The WF/PC-2 was used to make observations of the galaxy NGC-5457 (aka, M-101) to help improve the estimate of the Hubble Constant. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 3 Successful : 3 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Notes: There were no significant problems reported. Yesterday, an abbreviated ESTR#2 diagnostics test was conducted. Science fill data, at 1.024 Mbps, were recorded on track 1 from tape position BOT to approximately 40%, and on track 2 from 40% to EOT. Full playbacks were executed on two occasions. Preliminary evaluations from the DCF show no bit slips. Further analysis, based on data extractions of motor current and up-to-speed flag parameters, is proceeding. Additional playbacks of the data are also planned. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-